OpenStudy (anonymous):

Cot(theta) + tan(theta) =csc(theta)sec(theta) Prove the identity Can someone please help me with this problem?

OpenStudy (anonymous):

\[\cot \theta=\frac{ \cos \theta }{ \sin \theta }\] \[\tan \theta=\frac{ \sin \theta }{ \cos \theta }\] take L.C.M. of L.H.S and simplify